# Integral representation of solutions to initial-boundary value problems in the framework of the hyperbolic heat equation

### v1

@@ 4 A numerical example
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (4.1b)
@@
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=\left.\gamma_{l}e\right|_{x=l}.$ (4.3)
@@
where the functions $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ are given by
$\boldsymbol{a}^{\prime}(k,t)$ $=\begin{pmatrix}-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)\\[4.0pt] \mathrm{i}k\beta\tilde{e}^{(2)}(0,t)-\omega_{1}\tilde{g}^{(2)}(t)\end{pmatrix},$
$\boldsymbol{b}^{\prime}(k,t)$ $=\begin{pmatrix}\sigma_{l}^{(2)}(-k)\tilde{e}^{(1)}(l,t)\\[4.0pt] -\sigma_{l}^{(1)}(-k)\tilde{e}^{(2)}(l,t)\end{pmatrix},$
the functions $\tilde{e}^{(1,2)}$ are given by
$\begin{pmatrix}\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)\\[3.0pt] \tilde{e}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{1}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[3.0pt] 2\end{pmatrix}\omega_{2}\tilde{g}^{(1)}(t),$ (4.5a)
$\begin{pmatrix}\mathrm{i}k\beta\tilde{e}^{(2)}(0,t)\\[3.0pt] \tilde{e}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{2}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[3.0pt] 2\end{pmatrix}\omega_{1}\tilde{g}^{(2)}(t),$ (4.5b)
@@
$\varDelta_{1}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k),$
$\varDelta_{2}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k).$
@@
$-\mathrm{i}k\beta\tilde{e}^{(1)}(0,t)+\omega_{2}\tilde{g}^{(1)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t),\quad\tilde{e}^{(1)}(l,t)=\frac{2}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t),$
$-\mathrm{i}k\beta\tilde{e}^{(2)}(0,t)+\omega_{1}\tilde{g}^{(2)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t),\quad\tilde{e}^{(2)}(l,t)=\frac{2}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t).$
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{\omega_{1}-\omega_{2}}\bigg{[}\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{1}^{\prime}(k)}\omega_{2}\tilde{g}^{(1)}(t)-\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{2}^{\prime}(k)}\omega_{1}\tilde{g}^{(2)}(t)\bigg{]}.$ (4.6)
@@
Fig. 1 shows solutions to the problem (4.1)–(4.3) obtained by using the formula (4.4). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$. Here we use dimensionless values. The solution is calculated on the right boundary, i. e., for $x=1$. The solution to the problem with $\gamma_{l}=0$ (the thermally insulating boundary) and the general heat flux $g(t)$ can easily be obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. This solution is given by Eq. (A.5) (we have checked that in this case the formulae (4.4) and (A.5) give the same numerical results). In Ref. [37] the series solution to the problem (4.1)–(4.3) with $\gamma_{l}=0$ was obtained for the heat flux of the form (4.7).
@@
Figure 1: The solutions to the problem (4.1)–(4.3) on the right boundary, i. e., for $x=1$ with the heat flux $g(t)$ given by Eq. (4.7). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$.
5 Conclusions
In this paper we have considered the IBVPs on the finite interval for the system of the energy balance equation and modified Fourier’s law (constitutive equation). This system is hyperbolic. The boundary conditions include various models of behavior of a physical system at the boundaries. The most interesting model is when the behavior on at least one of the boundaries is described by Newton’s law, which states that the heat flux at the boundary is directly proportional to the difference in the temperature of the physical system and ambient temperature. In this case the boundary conditions express the relationship of unknown functions (temperature or internal energy and heat flux) with each other.
To solve the problems, we have applied the unified transform method. We did not solve problems in which one of the boundary conditions determines the temperature. These problems can be solved in the same way as in this paper. To illustrate the final formulae, we have considered a numerical example for a special case in which heat exchange at one of the boundaries obeys Newton’s law.
Also of considerable interest is the integral representation of solutions to IBVPs in the framework of the Guyer–Krumhansl model of heat conduction [41, 4, 5, 17, 42]. In the papers [37, 38, 43, 44, 45, 40] the IBVPs in the framework of the Guyer–Krumhansl model were solved both analytically and numerically for particular boundary conditions. The integral representation of the solutions to the initial value problems in the framework of the Guyer–Krumhansl model will be considered elsewhere.
@@ 5 Conclusions
Appendix A Particular cases that allow series representation of the solutions
In this appendix, for simplicity, we assume that heat sources are absent, i. e., $f=0$.
@@ Appendix A Particular cases that allow series representation of the solutions
A.1 The initial state is zero, the heat flux is prescribed on the left boundary, and the right boundary is thermally insulating
@@
Here we assume that the initial state is zero ($\varphi=0$, $\psi=0$), the heat flux is prescribed on the left boundary ($\gamma_{0}=0$), and the right boundary is thermally insulating ($\gamma_{l}=0$ and $h=0$). In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (A.1a)
$\tau\partial_{t}q+q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (A.1b)
@@
$\left.e\right|_{t=0}=0,\quad\left.q\right|_{t=0}=0,$ (A.2)
@@
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=0.$ (A.3)
@@
In this case the solution to the problem (A.1)–(A.3) is given by
$\boldsymbol{u}(x,t)=-\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (A.4)
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{(\omega_{1}-\omega_{2})(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})}\bigg{[}\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\omega_{1}\tilde{g}^{(2)}(t)-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\omega_{2}\tilde{g}^{(1)}(t)\bigg{]}.$
In order to obtain a series representation, the paths of integration $\partial D_{0}^{\pm}$ must be deformed back to the real axis. Note that the branch points $k=\pm(2\sqrt{\alpha\tau})^{-1}$ are removable. By deforming the paths, avoiding poles located on the real axis, and taking into account opposite directions of integration, we obtain contours around the poles. As a result of integration along the contours, the representation (A.4) is transformed into the series
@@
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$, $[(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ and $\omega_{1}(k_{n})\omega_{2}(k_{n})=\beta k_{n}^{2}$, we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{\mathrm{i}l(\omega_{1}^{(n)}-\omega_{2}^{(n)})}\begin{pmatrix}\omega_{1}^{(n)}\tilde{g}^{(2)}(t)-\omega_{2}^{(n)}\tilde{g}^{(1)}(t)\\[5.0pt] -\mathrm{i}k_{n}\beta\left[\tilde{g}^{(2)}(t)-\tilde{g}^{(1)}(t)\right]\end{pmatrix},\quad n\neq 0,$
@@
where $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$, and
@@
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}\\[2.0pt] =\frac{1}{l}\begin{pmatrix}\tilde{g}(t)\\[3.0pt] 0\end{pmatrix}+\frac{2}{l}\sum_{n=1}^{\infty}\frac{1}{\omega_{1}^{(n)}-\omega_{2}^{(n)}}\begin{pmatrix}\left[\omega_{1}^{(n)}\tilde{g}^{(2)}(t)-\omega_{2}^{(n)}\tilde{g}^{(1)}(t)\right]\cos(k_{n}x)\\[3.0pt] k_{n}\beta\left[\tilde{g}^{(2)}(t)-\tilde{g}^{(1)}(t)\right]\sin(k_{n}x)\end{pmatrix},$ (A.5)
@@
where $k_{n}=\pi n/l$, $\omega_{1,2}^{(n)}\equiv\omega_{1,2}(k_{n})$ (see Eq. (3.1)), $\tilde{g}^{(1,2)}(t)$ are given by Eq. (3.19a). This solution can also be easily obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. In Ref. [37] the solution was obtained for a particular heat flux $g(t)$ of the form (4.7).
A.2 The initial state is non-zero and the boundaries are thermally insulating

### v2

@@ 4 A numerical example
$\tau\partial_{t}q+q-\mu^{2}\partial_{x}^{2}q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (4.1b)
@@
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=\left.\gamma_{l}e\right|_{x=l}$ (4.3)
(the second boundary condition expresses Newton’s law, when the ambient temperature is zero).
@@
where the vectors $\boldsymbol{a}^{\prime}$ and $\boldsymbol{b}^{\prime}$ are given by
$\boldsymbol{a}^{\prime}(k,t)$ $=\begin{pmatrix}-\mathrm{i}k(\beta-\theta\omega_{1})\tilde{e}_{0}^{(1)}(0,t)+(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)\\[3.0pt] \mathrm{i}k(\beta-\theta\omega_{2})\tilde{e}_{0}^{(2)}(0,t)-(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)\end{pmatrix}$,
$\boldsymbol{b}^{\prime}(k,t)$ $=\begin{pmatrix}\sigma_{l}^{(1)}(-k)\tilde{e}_{0}^{(1)}(l,t)\\[3.0pt] -\sigma_{l}^{(2)}(-k)\tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}$,
(see Eq. (3.34)), the coefficients $\sigma_{l}^{(1,2)}(k)$ are given by Eq. (3.26), the functions $\tilde{e}_{0}^{(1,2)}$ are given by
$\begin{pmatrix}\mathrm{i}k(\beta-\theta\omega_{1})\tilde{e}_{0}^{(1)}(0,t)\\[3.0pt] \tilde{e}_{0}^{(1)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{1}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k)\\[3.0pt] 2\end{pmatrix}(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t),$ (4.5a)
$\begin{pmatrix}\mathrm{i}k(\beta-\theta\omega_{2})\tilde{e}_{0}^{(2)}(0,t)\\[3.0pt] \tilde{e}_{0}^{(2)}(l,t)\end{pmatrix}$ $=\frac{1}{\varDelta_{2}^{\prime}(k)}\begin{pmatrix}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k)\\[3.0pt] 2\end{pmatrix}(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t),$ (4.5b)
@@
$\begin{split}\varDelta_{1}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(1)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(1)}(-k),\\ \varDelta_{2}^{\prime}(k)=\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}\sigma_{l}^{(2)}(k)+\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\sigma_{l}^{(2)}(-k).\end{split}$
@@
$\begin{split}-\mathrm{i}k(\beta-\theta\omega_{1})\tilde{e}_{0}^{(1)}(0,t)+(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{1}^{\prime}(k)}(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t),\\ -\mathrm{i}k(\beta-\theta\omega_{2})\tilde{e}_{0}^{(2)}(0,t)+(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)=2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{2}^{\prime}(k)}(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t).\end{split}$
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{\omega_{1}-\omega_{2}}\bigg[\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(1)}(-k)}{\varDelta_{1}^{\prime}(k)}(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)-\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}\frac{\sigma_{l}^{(2)}(-k)}{\varDelta_{2}^{\prime}(k)}(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)\bigg].$ (4.6)
@@
Figs. 1, 2 show solutions to the problem (4.1)–(4.3) obtained by using the formula (4.4). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\mu^{2}=0.02$ and $\mu^{2}=0.2$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$. Here we use dimensionless values. The solution is calculated on the right boundary, i. e., for $x=1$. The solution to the problem with $\gamma_{l}=0$ (the thermally insulating boundary) and the general heat flux $g(t)$ can easily be obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. This solution is given by Eq. (B.5) (we have checked that in this case the formulae (4.4) and (B.5) give the same numerical results). In Ref. [6] the series solution to the problem (4.1)–(4.3) with $\gamma_{l}=0$ was obtained for the heat flux of the form (4.7).
@@
Figure 1: The solutions to the problem (4.1)–(4.3) on the right boundary, i. e., for $x=1$ with the heat flux $g(t)$ given by Eq. (4.7). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\mu^{2}=0.02$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$.
Figure 2: The solutions to the problem (4.1)–(4.3) on the right boundary, i. e., for $x=1$ with the heat flux $g(t)$ given by Eq. (4.7). The values of the parameters are $\alpha=1$, $\tau=0.02$, $\mu^{2}=0.2$, $\gamma_{l}=0$ and $\gamma_{l}=0.2$, $l=1$, $\tau_{\Delta}=0.04$.
Appendix A The paths of integration in the complex plane
To find the paths of integration we must solve the equation
$\operatorname{Re}F(k)=0,\quad\text{where}\quad F(k)=1+\mu^{2}k^{2}\pm\sqrt{(1+\mu^{2}k^{2})^{2}-4\alpha\tau k^{2}},$
see Eq. (3.1). Making the change of variable $\mu k=z\equiv x+\mathrm{i}y$ ($x=\operatorname{Re}z$, it is not the spatial variable, as in the main part of the paper), and introducing the parameter $C=\alpha\tau/\mu^{2}$, we obtain
$F(k)\equiv G(z)=1+z^{2}\pm\sqrt{(1+z^{2})^{2}-4Cz^{2}}\\ =x^{2}-y^{2}+1+\mathrm{i}2xy\pm\sqrt{(x^{2}-y^{2}+1+\mathrm{i}2xy)^{2}-4C(x^{2}-y^{2}+\mathrm{i}2xy)}\\ =x^{2}-y^{2}+1+\mathrm{i}2xy\pm\sqrt{\xi+\mathrm{i}\eta}\\ =x^{2}-y^{2}+1+\mathrm{i}2xy\pm\sqrt{r^{2}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}2\varphi}}=x^{2}-y^{2}+1+\mathrm{i}2xy\pm r\mathop{\mathrm{e}}\nolimits^{\mathrm{i}\varphi},$
where
$r^{2}=\sqrt{\xi^{2}+\eta^{2}},\quad\cot 2\varphi=\frac{\xi}{\eta},\quad\varphi\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right),$
$\xi=(x^{2}-y^{2}+1)^{2}-4x^{2}y^{2}-4C(x^{2}-y^{2}),\quad\eta=4xy(x^{2}-y^{2}+1-2C).$
So the equation becomes
$\operatorname{Re}G(z)=0\quad\Leftrightarrow\quad r\cos\varphi=|x^{2}-y^{2}+1|\quad\Leftrightarrow\quad r^{2}\cos^{2}\varphi=(x^{2}-y^{2}+1)^{2}.$
Using the equality
$\cos^{2}\varphi=\frac{1}{2}\left(1+\operatorname{sign}\varphi\frac{\cot 2\varphi}{\sqrt{\cot^{2}2\varphi+1}}\right),\quad\varphi\in\left(-\frac{\pi}{2},\frac{\pi}{2}\right),$
we obtain the equation
$\frac{r^{2}}{2}\left(1+\operatorname{sign}\varphi\frac{\xi}{\eta\sqrt{(\xi/\eta)^{2}+1}}\right)\equiv\frac{r^{2}}{2}\left(1+\frac{\xi}{\sqrt{\xi^{2}+\eta^{2}}}\right)\equiv\frac{\sqrt{\xi^{2}+\eta^{2}}+\xi}{2}\\ =(x^{2}-y^{2}+1)^{2},$
which is equavalent to the equation
$\xi^{2}+\eta^{2}=\left[2(x^{2}-y^{2}+1)^{2}-\xi\right]^{2}.$
As a result, we get the equation
$y^{6}+(x^{2}-2)y^{4}-(x^{4}-4Cx^{2}-1)y^{2}-(x^{6}+2x^{4}+x^{2})=0$ (A.1)
that is actually a cubic equation for $y^{2}$. This equation can be written in the form
$(y^{2}-x^{2})[(x^{2}+y^{2}-1)^{2}+4x^{2}]+4(C-1)x^{2}y^{2}=0,$
or, equivalently
$y=\pm\sqrt{\dfrac{(x^{2}+y^{2}-1)^{2}+4x^{2}}{(x^{2}+y^{2}-1)^{2}+4Cx^{2}}}|x|,\quad C=\frac{\alpha\tau}{\mu^{2}}.$ (A.2)
This relation expresses the dependence of $y$ on $x$ implicitly, however, it allows us to obtain qualitative information about the solutions of the equation (A.1). The relation (A.2) shows that the equation (A.1) has exactly two real solutions $y(x)$, and, therefore, the paths $\partial D^{\pm}$, Eq. (3.14), are uniquely determined. If $C=1$, then the paths $\partial D^{\pm}$ are determined by the equations $k_{\mathrm{I}}=\pm|k_{\mathrm{R}}|$, respectively. Note that these paths coincide with the paths of integration for the heat equation on a finite interval, see Ref. [24]. If $C<1$, then the path $\partial D^{+}$ passes above the path $k_{\mathrm{I}}=|k_{\mathrm{R}}|$, and the path $\partial D^{-}$ passes below the path $k_{\mathrm{I}}=-|k_{\mathrm{R}}|$. If $C>1$, then the path $\partial D^{+}$ passes below the path $k_{\mathrm{I}}=|k_{\mathrm{R}}|$ and above the real axis, and the path $\partial D^{-}$ passes above the path $k_{\mathrm{I}}=-|k_{\mathrm{R}}|$ and below the real axis. The relation (A.2) shows also the asymptotic behavior
$\left|\frac{y}{x}\right|\to 1\quad\text{as}\quad x\to 0\enskip\text{or}\enskip x\to\infty.$
To determine the paths $\partial D^{\pm}$ explicitly, we write the equation (A.1) in the form
$s^{3}+as^{2}+bs+c=u^{3}+Pu+Q=(u-2A)[(u+A)^{2}+B^{2}]=0,$ (A.3)
where $y^{2}\equiv s=u-a/3$,
$a=x^{2}-2,\quad b=-(x^{4}-4Cx^{2}-1),\quad c=-(x^{6}+2x^{4}+x^{2}),$
$P=-\frac{a^{2}}{3}+b,\quad Q=2\left(\frac{a}{3}\right)^{3}-\frac{a}{3}b+c,\quad A=\frac{\alpha+\beta}{2},\quad B=\sqrt{3}\,\frac{\alpha-\beta}{2},$
$\alpha=\sqrt[3]{-\frac{Q}{2}+\sqrt{D}},\quad\beta=\sqrt[3]{-\frac{Q}{2}-\sqrt{D}},\quad D=\left(\frac{P}{3}\right)^{3}+\left(\frac{Q}{2}\right)^{2},$
the roots $\alpha$ and $\beta$ are chosen so that the equality $\alpha\beta=-P/3$ is valid and the value AA is real. The real nonnegative root of the equation (A.3) is given by
$y^{2}\equiv s=2A-\frac{a}{3}.$
This relation determines the paths $\partial D^{\pm}$.
@@ Appendix A The paths of integration in the complex plane
Appendix B Particular cases that allow series representation of the solutions
@@
B.1 The heat sources are absent, the initial state is zero, the heat flux is prescribed on the left boundary, and the right boundary is thermally insulating
@@
Here we assume that heat sources are absent ($f=0$), the initial state is zero ($\varphi=0$, $\psi=0$), the heat flux is prescribed on the left boundary ($\gamma_{0}=0$), and the right boundary is thermally insulating ($\gamma_{l}=0$ and $h=0$). In this case both equations are homogeneous:
$\partial_{t}e+\partial_{x}q=0,\quad x\in(0,l),\quad t>0,$ (B.1a)
$\tau\partial_{t}q+q-\mu^{2}\partial_{x}^{2}q+\alpha\partial_{x}e=0,\quad x\in(0,l),\quad t>0.$ (B.1b)
@@
$\left.e\right|_{t=0}=0,\quad\left.q\right|_{t=0}=0,$ (B.2)
@@
$\left.q\right|_{x=0}=g(t),\quad\left.q\right|_{x=l}=0.$ (B.3)
@@
In this case the solution to the problem (B.1)–(B.3) is given by
$\boldsymbol{u}(x,t)=-\frac{1}{2\pi}\int_{\partial D_{0}^{+}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\mathop{}\!\mathrm{d}k-\frac{1}{2\pi}\int_{\partial D_{0}^{-}}\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\mathop{}\!\mathrm{d}k,$ (B.4)
@@
$S(k)\boldsymbol{a}^{\prime}(k,t)=\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\\[2.0pt] =\frac{2\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}}{(\omega_{1}-\omega_{2})(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})}\bigg[\begin{pmatrix}1\\ \omega_{2}/\mathrm{i}k\end{pmatrix}(\omega_{1}-\theta k^{2})\tilde{g}^{(2)}(t)-\begin{pmatrix}1\\ \omega_{1}/\mathrm{i}k\end{pmatrix}(\omega_{2}-\theta k^{2})\tilde{g}^{(1)}(t)\bigg]$
(see Eq. (4.6)).
In order to obtain a series representation, the paths of integration $\partial D_{0}^{\pm}$ must be deformed back to the real axis. Note that the branch points $k=(\pm\sqrt{\alpha\tau}\pm\sqrt{\alpha\tau-\mu^{2}})/\mu^{2}$ are removable. By deforming the paths, avoiding poles located on the real axis, and taking into account opposite directions of integration, we obtain contours around the poles. As a result of integration along the contours, the representation (B.4) is transformed into the series
@@
Using the equalities $\mathop{\mathrm{e}}\nolimits^{\pm\mathrm{i}k_{n}l}=(-1)^{n}$, $(\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kl}-\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl})_{k}^{\prime}]|_{k=k_{n}}=2\mathrm{i}l(-1)^{n}$ and $\omega_{1}(k_{n})\omega_{2}(k_{n})=\beta k_{n}^{2}$, we see that the residues are equal to
$\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}S(k)\boldsymbol{a}^{\prime}(k,t)\right]=\operatorname*{Res}_{k=k_{n}}\left[\mathop{\mathrm{e}}\nolimits^{\mathrm{i}kx}\mathop{\mathrm{e}}\nolimits^{-\mathrm{i}kl}S(k)\boldsymbol{b}^{\prime}(k,t)\right]\\[2.0pt] =\frac{\mathop{\mathrm{e}}\nolimits^{\mathrm{i}k_{n}x}}{\mathrm{i}l(\omega_{1,n}-\omega_{2,n})}\begin{pmatrix}(\omega_{1,n}-\theta k_{n}^{2})\tilde{g}^{(2)}(t)-(\omega_{2,n}-\theta k_{n}^{2})\tilde{g}^{(1)}(t)\\[5.0pt] -\mathrm{i}k_{n}\big[(\beta-\theta\omega_{2,n})\tilde{g}^{(2)}(t)-(\beta-\theta\omega_{1,n})\tilde{g}^{(1)}(t)\big]\end{pmatrix},\quad n\neq 0,$
@@
where $\omega_{m,n}\equiv\omega_{m}(k_{n})$, and
@@
$\boldsymbol{u}(x,t)\equiv\begin{pmatrix}e(x,t)\\[2.0pt] q(x,t)\end{pmatrix}=\frac{1}{l}\begin{pmatrix}\tilde{g}(t)\\[3.0pt] 0\end{pmatrix}\\[2.0pt] +\frac{2}{l}\sum_{n=1}^{\infty}\frac{1}{\omega_{1,n}-\omega_{2,n}}\begin{pmatrix}\big[(\omega_{1,n}-\theta k_{n}^{2})\tilde{g}^{(2)}(t)-(\omega_{2,n}-\theta k_{n}^{2})\tilde{g}^{(1)}(t)\big]\cos(k_{n}x)\\[3.0pt] k_{n}\left[(\beta-\theta\omega_{2,n})\tilde{g}^{(2)}(t)-(\beta-\theta\omega_{1,n})\tilde{g}^{(1)}(t)\right]\sin(k_{n}x)\end{pmatrix}$, (B.5)
@@
where $\beta=\alpha/\tau$, $\theta=\mu^{2}/\tau$, $k_{n}=\pi n/l$, $\omega_{m,n}\equiv\omega_{m}(k_{n})$ (see Eq. (3.1)), $\tilde{g}^{(1,2)}(t)$ are given by Eq. (3.20a). This solution can also be easily obtained by the Fourier method with preliminary elimination of inhomogeneities in boundary conditions. In Ref. [6] the solution was obtained for a particular heat flux $g(t)$ of the form (4.7).
B.2 The initial state is non-zero and the boundaries are thermally insulating